# MiMiC: A High-Performance Framework for Multiscale Molecular Dynamics Simulations

### v1

@@ III Workflow Examples
The overall workflow for multiscale simulations varies depending on the system under study and the specific multiscale method employed. For multiscale QM/MM MD simulations, we typically start with a system that has been equilibrated at the classical MM level. This MM equilibration phase involves the usual steps for classical MD simulations, such as energy minimization and equilibration to target conditions (e.g., ambient temperature) in the selected thermodynamic ensemble. Subsequently, input files for all programs involved in the simulation are prepared. This includes providing MiMiC with the information needed for client communication, the definition of subsystems, and the computation of subsystem interactions. The MiMiCPy preprocessing toolkit, described in Section II.4, is designed to facilitate this part of the workflowRaghavan et al. (2023a). Following the preparation, the multiscale simulation is performed. Like in any classical or ab inito MD workflow, this includes benchmarking to ensure optimal use of computational resources, equilibration, and production simulations. The generated trajectory data can be post-processed to evaluate various properties of interest. Although we currently rely on the output files produced by the external programs that can be analyzed using standard tools for MD analysis, we plan to introduce our own output to facilitate and homogenize the post-processing phase.
MiMiC-based multiscale simulations can proceed in several different ways depending on the method and its implementation. In this section, we describe in detail three such workflows: electrostatic embedding QM/MM in Section III.1, polarizable embedding QM/MM in Section III.2, and MTS QM/MM in Section III.3.
@@ III.1 Electrostatic Embedding QM/MM MD
In electrostatic embedding QM/MM, the fixed point charges (or multipoles) of the MM subsystem generate an external field that polarizes the electron density of the QM subsystem. The QM equations are thus solved in the presence of this field, but since the MM point charges are fixed, the MM subsystem is not explicitly polarized. A possible way of implementing QM/MM is to employ an additive scheme, in which the total time-independent Hamiltonian of the system,
@@
Before proceeding to the treatment of the long-range interactions, we introduce a multi-index notation, which allows us to express the equations involving multipoles in a more compact manner. A multi-index $\alpha$ is a tuple consisting of three non-negative integers associated with the Cartesian components, i.e. $\alpha=(\alpha_{x},\alpha_{y},\alpha_{z})$. The multi-index norm is defined as $\left|\alpha\right|=\alpha_{x}+\alpha_{y}+\alpha_{z}$, the multi-index factorial as $\alpha!=\alpha_{x}!\cdot\alpha_{y}!\cdot\alpha_{z}!$, and the multi-index partial derivative operator as $\partial^{\alpha}_{\mathbf{R}_{b}}=\partial^{\left|\alpha\right|}/\partial^{\alpha_{x}}_{R_{b,x}}\partial^{\alpha_{y}}_{R_{b,y}}\partial^{\alpha_{z}}_{R_{b,z}}$. Using this notation, the $\alpha$-component of a tensor is denoted using square brackets as, e.g., $T^{[\alpha]}$, while a tensor is denoted with its rank in round parentheses, e.g., $T^{(0)}$.
@@
where $\alpha$ is a multi-index indicating the Cartesian component and $\bar{\mathbf{R}}_{\mathrm{QM}}$ is the origin of the expansion, e.g., the centroid of the QM subsystem $\sum_{j=1}^{N_{\mathrm{QM}}}\mathbf{R}_{\mathrm{QM},j}/N_{\mathrm{QM}}$. The electrostatic energy for the long-range interactions can then be expressed as
@@
$T^{[\alpha]}(\mathbf{R}_{a},\mathbf{R}_{b})=\partial^{\alpha}_{\mathbf{R}_{b}}\frac{1}{\left|\mathbf{R}_{b}-\mathbf{R}_{a}\right|}\ .$ (8)
This approach drastically reduces the overall computational cost by increasing the accuracy of the long-range contributions, allowing the selection of smaller short-range regions, whose interactions are substantially more expensive to compute. In particular, with a proper selection of the short-range cutoff, which dictates how many MM atoms are treated in short- and long-range domains, and the order for the multipole expansion, it is possible to reach the same accuracy as if all the electrostatic interactions were treated exactly, i.e. according to Eq. III.1, yet, at a fraction of the costOlsen et al. (2019).
A detailed theoretical treatment of the electrostatic embedding QM/MM scheme implemented using MiMiC can be found in Ref. 33. It includes explicit expressions for the polarizing potential acting on the QM subsystem due to MM atoms, i.e., the functional derivative of the QM/MM interaction energy with respect to the electron density, and for the forces acting on QM and MM atoms, i.e., the derivatives of the QM/MM interaction energy with respect to the atomic positions.
@@
Electrostatic embedding QM/MM is implemented using MiMiC by interfacing two external programs, a QM and an MM client, to an MD driver. The general simulation workflow is illustrated in Figure 4. After the initialization, consisting of data collection and allocation, the simulation enters the main MD loop. At each MD step, MiMiC distributes atomic coordinates among the clients, which then concurrently compute energy and forces that are collected by MiMiC and passed to the MD driver to update velocities and subsequently calculate a new set of coordinates.
@@
The QM client solves the QM equations in the presence of the external MM potential computed by MiMiC. The client then sends the converged electron density to MiMiC, and finally proceeds to compute the QM forces, while MiMiC computes the corresponding electrostatic QM/MM contributions. Finally, MiMiC collects energies and forces from all client programs and forwards them to the MD driver, which then updates velocities and positions, and starts a new MD step.
@@ III.2 Polarizable Embedding QM/MM MD
In electrostatic embedding, the QM system is polarized by the electric field generated by the fixed point charges of the MM subsystem. As such, this approach does not account for the explicit polarization of the MM subsystem, which is important in many cases and even crucial in some, e.g., in processes involving absorption or emission of photons and electron or proton transferBondanza et al. (2020). Polarizable embedding QM/MM schemes address this issue by using an MM force field that includes a polarizable component. For instance, in the frequently used induced point dipole model, the additional dynamic degrees of freedom are modeled using electric polarizabilities that describe an electronic charge distribution distorted by an electric field. Here, we briefly summarize the theoretical basis for polarizable embedding QM/MM MD using the polarizable AMOEBAPonder et al. (2010) force field to model the MM subsystem (QM/AMOEBA). For further details, we refer to Refs. 45; 46.
The electrostatic interactions in the AMOEBA force field are modeled using permanent multipoles, up to and including quadrupoles, and induced dipoles, from which electrostatic and polarization QM/MM interaction energies arise. The former is computed similarly as in electrostatic embedding QM/MM using modified interaction tensors (Eq. 5). The polarization energy is expressed in terms of induced dipoles and electric fields, thus making it necessary to first determine the induced dipoles through coupled polarization equations. A characteristic aspect of AMOEBA is its non-variational nature stemming from the fact that the electric field from the permanent multipoles used to determine the induced dipoles differs from the one used to calculate the polarization energy. The former is referred to as the direct field and the latter as the polarization field. The difference appears due to the use of different scaling factors and exclusion rules for each of these fields. Consequently, it is not possible to derive the polarization equations by simply requiring the stationarity of the energy functional with respect to the polarizable degrees of freedom. To avoid the complexity of non-variational optimization, a variational formulation has been adopted, resulting in two sets of polarization equations and, therefore, two sets of induced dipolesLoco et al. (2017); Nottoli and Lipparini (2020).
@@
Enabling fully polarizable QM/MM MD simulations using MiMiC requires modifications of the algorithm used to solve the QM equations. Achieving full self-consistency entails that the induced dipoles in the MM subsystem and the electron density in the QM subsystem are iteratively adjusted until both are converged with respect to the field generated by the other subsystem. This iterative process is characteristic of most polarizable embedding QM/MM models. The general workflow of a polarizable embedding QM/MM MD simulation using MiMiC is quite similar to the one for electrostatic embedding QM/MM shown in Figure 4. It requires extending the Solve QM equations step with the iterative cycle illustrated in Figure 5.
@@ III.3 Multiple Time Step QM/MM MD
MTS algorithms reduce the cost of MD simulations by separating forces into fast and slow components, and integrating the latter less frequently, i.e., with a larger time step. Assuming a time-scale separation between these components, it is possible to split the classical Liouville operator of an $N$ particle system as
@@
MTS-generated trajectories are, by construction, at the high level of theory with outer time step $\Delta t$, regardless of the low-level method used. The choice of the latter only determines the magnitude of the correction and hence the oscillations of the slow force component over time, which dictate the maximal ratio between the inner and the outer steps that can be applied to generate stable dynamics.
@@
Performing a full MTS MD step requires executing $n$ inner steps at a lower level of theory (as shown in Eq. 12) followed by a single high-level outer step. Figure 6 depicts a schematic workflow of an MTS QM/MM MD algorithm using MiMiC. At every step, MiMiC collects the low-level forces from the QM and MM clients and computes the low-level QM/MM force contributions (Eq. 14a). If the number of steps is not a multiple of $n$, velocities and coordinates are updated at the low level with a time step $\Delta t/n$, otherwise, at every $n$-th step, the outer loop is undertaken. In the outer loop, MiMiC additionally collects the high-level forces and computes the high-level QM/MM force contributions (Eq. 14b), and proceeds to calculate the slow forces. These are used by the MD driver to update the velocities, thus completing a full MD cycle with the time step $\Delta t$ at the high level of theory. Note that with MiMiC, it is straightforward to use two different QM client programs for the low- and high-level forces.
IV Features and Client Programs
In this section, we briefly discuss the present and upcoming features of MiMiC. As multiscale QM/MM simulations remain the focus of our immediate development plans, we place particular emphasis on new QM and MM client programs that are being coupled to the framework.
The first development version of the MiMiC framework was released in 2022 on the occasion of the CECAM Flagship School Multiscale Molecular Dynamics with MiMiCmim (2022) held at the CECAM Headquarters at EPFL in Lausanne, Switzerland. It implements the DFT-based electrostatic embedding QM/MM MD method illustrated in Subsection III.1, using GROMACS Abraham et al. (2015); Páll et al. (2015) as the MM client and CPMD cpm (2023) as the QM client and MD driver. Furthermore, it also supports MTS QM/MM MD (Subsection III.3), and, through the PLUMED library, QM/MM-based enhanced sampling simulations and free energy calculationsBonomi et al. (2009); Tribello et al. (2014); Colón-Ramos et al. (2019).
@@ IV Features and Client Programs
Our initial choice for an MM client was GROMACSPáll et al. (2015), which is a free and open-source software suite for classical MD, distributed under LGPL version 2.1, whose developers aimed to provide the highest possible performance and efficiency on any hardwareAbraham et al. (2015). It implements a native heterogeneous parallelization setup, using both CPUs and GPUs, with the possibility to offload all force components to GPUs in single-precision calculations. The latest versionsPáll et al. (2020) also introduced new direct GPU–GPU communication and extended GPU integration, enabling excellent performance across multiple GPUs and efficient multi-node parallelization. Moreover, GROMACS supports a large variety of popular biomolecular force fields, including AMBERPonder and Case (2003), CHARMMBrooks et al. (2009), GROMOSScott et al. (1999), and OPLSRobertson, Tirado-Rives, and Jorgensen (2015), and provides a wide array of well-documented pre- and post-processing tools. New users can also benefit from an extensive range of tutorials, documentation, and an active user community.
Figure 7: Parallel scalability of a MiMiC-based QM/MM implementation using GROMACS as the MM client and CPMD as the QM client and MD driver. The system comprised 130 828 atoms, out of which 142 were treated at the QM level using the hybrid B3LYP functional. The simulations were run on the CPU partition of the JUWELS cluster at the Jülich Supercomputing Centre. Adapted from Raghavan et al.Raghavan et al. (2023b), licensed under CC BY 4.0.
@@
IV.1 Upcoming Features
@@
These efforts include the implementation of dynamically generated restrained electrostatic potential (D-RESP) chargesLaio, VandeVondele, and Rothlisberger (2002b), extending the original approach to also fit higher-order atom-centered multipoles. Such an extension will also make it possible to implement different coupling schemes for electrostatic interactions, for example, by introducing a three-level approachLaio et al. (2004), which would significantly increase electrostatic embedding efficiency, particularly, for elongated QM subsystems. In addition, the implementation of D-RESP charges will also serve as an on-the-fly analysis tool for tracking ongoing changes in the electronic structure and will also enable charge-driven enhanced samplingSulpizi, Rothlisberger, and Laio (2005). Another important addition, building on D-RESP, will be the development of QM/MM force-matching schemes, such as those in Refs. 89 and 90, in which the total forces on the QM atoms, as well as the QM electrostatic potential and field, are extracted during a QM/MM MD simulation, for a set of sampled configurations. These quantities are then used to parameterize standard or polarizable biophysical force fields for the QM subsystem. Such QM-matched force fields can then be employed to achieve more extensive sampling in a fully classical MD simulation of the whole system.
@@
A different extension of capabilities in MiMiC will be based on enhancing the MTS workflow illustrated in Section III.3, which unlocks a plethora of possible combinations of different high- and low-level methods, including wave function, DFT, semiempirical, force field, and ML modelsMouvet et al. (2022). A future release of MiMiC will also include an adaptive ML-accelerated MTS QM/MM scheme where the ML model is retrained on the fly whenever the system accesses new regions of configurational space (e.g. during a chemical reaction) for which the original ML model has insufficient predictive power. In addition, this adaptive scheme allows for a dynamic training set build-up during the run, thus, circumventing the need for extensive and computationally demanding a priori generation of high-level reference data.
@@
Concerning the current developments of new interfaces to external client programs, we now primarily focus on the following three objectives. First, we strive to enable large-scale QM/MM MD simulations on current and future (pre-)exascale supercomputers by coupling to highly scalable GPU-enabled QM client programs, namely, CP2K Kühne et al. (2020), DFT-FE Motamarri et al. (2020); Das et al. (2022), and Quantum ESPRESSO Giannozzi et al. (2009); Carnimeo et al. (2023), as well as making available more accurate wave function methods for QM/MM, as implemented in CP2K and CFOUR Matthews et al. (2020); Stanton et al. . Finally, we aim to implement high-performance, GPU-accelerated polarizable embedding QM/MM (see Section III.2) via polarizable force fields as implemented in Tinker-HP Lagardère et al. (2018); Adjoua et al. (2021) and OpenMM Eastman et al. (2012, 2017, 2023).
@@
To achieve these goals, work on interfaces with several client programs is currently underway. In the remainder of this section, we review the distinctive features that make them ideal candidates for integration into the framework. As new programs will be coupled to MiMiC, the pre-processing toolkit MiMiCPy will also be augmented accordingly.
@@
CP2KKühne et al. (2020) is a free and open-source software package, distributed under GNU General Public License (GPL) version 2.0 or later, for simulations of solid-state and (bio)molecular systems, giving access to a broad range of electronic structure methods. It focuses on ensuring excellent performance, thanks to algorithms developed for modern HPC systems, also including GPU support. In particular, the electronic structure module QuickstepKrack and Parrinello (2004); VandeVondele et al. (2005); Kühne et al. (2020) provides an efficient infrastructure of integral routines and optimization algorithms for a wide spectrum of methods, including orbital-free DFT, KS-DFT, and MP2. Quickstep is based on mixed Gaussian and PW (GPW) basis sets and their augmented extension (GAPW), where auxiliary PW basis sets are used within a GTO scheme. Recently, CP2K’s linear scaling DFT implementation was used for calculations on the SARS-CoV-2 spike proteins with almost 83 million atoms, for which it scaled up to 4 400 NVIDIA A100 GPUs with a parallel efficiency of over 80%Schade et al. (2023).
@@
DFT-FE Motamarri et al. (2020); Das et al. (2022) is a recently developed free and open-source code, distributed under LGPL version 2.1 or later, that implements real-space KS-DFT based on a spatially adaptive and systematically convergent higher-order spectral finite-element (FE) basis. It accommodates both norm-conserving pseudopotentials and all-electron calculations, with ongoing implementation efforts to support ultrasoft projector-augmented-wave pseudopotentials. DFT-FE enables fast and accurate large-scale DFT calculations on generic materials systems involving many tens of thousands of electrons on both many-core CPU and hybrid CPU-GPU architectures. The good parallel scalability and efficient use of GPUs has enabled systematically convergent DFT calculations with low wall times, e.g., an ab initio MD step of systems containing 10 000-20 000 electrons can be completed in wall times of <1 minute Das et al. (2022). In addition to the support of NVIDIA GPUs through CUDA, DFT-FE’s GPU porting layer was recently extended to support AMD GPUs through HIP, with its exascale capability recently demonstrated through scaling up to 8 000 GPU nodes (32 000 AMD MI250X GPUs) of the Oak Ridge Leadership Computing Facility Frontier exascale supercomputer Das et al. (2023).
@@
Quantum ESPRESSOGiannozzi et al. (2020) is an integrated and free suite of programs, distributed under GPL version 2.0 or later, designed for first-principles electronic structure calculations and materials modeling at the nanoscale. In recent years, it has been refactored and organized in multiple layers of codesCarnimeo et al. (2023), consisting of loosely coupled modules and libraries, tailored to be independently extensible and maintainable. Recent versions are composed of a core computational layer for DFT simulations, based on PWs and pseudopotentials, containing a quantum-engine package to solve self-consistent KS equations for periodic systems (PWscf). To achieve high performance on different architectures, Quantum ESPRESSO features a layered, fine-tuneable parallelization scheme based on MPI and OpenMP, but also GPU support. Porting to GPU has been achieved through CUDA Fortran and, to an increasing degree, directive-based programming models (OpenACC and OpenMP) for offloading to heterogeneous platforms, making it as architecture-agnostic as possible. The optimal combination of all available parallelism levels currently allows the Quantum ESPRESSO suite to achieve a substantial increase in performance and efficiencyCarnimeo et al. (2023).

### v2

@@ III Workflow Examples
The overall workflow for multiscale simulations varies depending on the system under study and the specific multiscale method employed. For multiscale QM/MM MD simulations, we typically start with a system that has been equilibrated at the classical MM level. This MM equilibration phase involves the usual steps for classical MD simulations, such as energy minimization and equilibration to target conditions (e.g., ambient temperature) in the selected thermodynamic ensemble. Subsequently, input files for all programs involved in the simulation are prepared. This includes providing MiMiC with the information needed for client communication, the definition of subsystems, and the computation of subsystem interactions. The MiMiCPy preprocessing toolkit, described in Sec. II.4, is designed to facilitate this part of the workflowRaghavan et al. (2023a). Following the preparation, the multiscale simulation is performed. Like in any classical or ab inito MD workflow, this includes benchmarking to ensure optimal use of computational resources, equilibration, and production simulations. The generated trajectory data can be post-processed to evaluate various properties of interest. Although we currently rely on the output files produced by the external programs that can be analyzed using standard tools for MD analysis, we plan to introduce our own output to facilitate and homogenize the post-processing phase.
MiMiC-based multiscale simulations can proceed in several different ways depending on the method and its implementation. In this section, we describe in detail three such workflows: electrostatic embedding QM/MM in Sec. III.1, polarizable embedding QM/MM in Sec. III.2, and MTS QM/MM in Sec. III.3.
@@ III.1 Electrostatic Embedding QM/MM MD
In electrostatic embedding QM/MM, the fixed point charges (or multipoles) of the MM subsystem generate an external field that polarizes the electron density of the QM subsystem. The QM equations are thus solved in the presence of this field, but since the MM point charges are fixed, the MM subsystem is not explicitly polarized. A possible way of implementing QM/MM is to employ an additive scheme in which the total time-independent Hamiltonian of the system,
@@
where $\bar{\mathbf{R}}_{\mathrm{QM}}$ is the origin of the expansion, usually the centroid of the QM subsystem, and $\alpha=(\alpha_{x},\alpha_{y},\alpha_{z})$ is a multi-indexRaymond (1991) indicating the Cartesian component. In this notation, the sum of components and factorial are defined as $\left|\alpha\right|=\alpha_{x}+\alpha_{y}+\alpha_{z}$ and $\alpha!=\alpha_{x}!\cdot\alpha_{y}!\cdot\alpha_{z}!$, respectively. For more details, we refer to Ref. 34. The electrostatic energy for the long-range interactions can then be expressed as
@@
$T^{[\alpha]}(\mathbf{R}_{a},\mathbf{R}_{b})=\frac{\partial^{\left|\alpha\right|}}{\partial R_{b,\,x}^{\alpha_{x}}\partial R_{b,\,y}^{\alpha_{y}}\partial R_{b,\,z}^{\alpha_{z}}}\frac{1}{\left|\mathbf{R}_{a}-\mathbf{R}_{b}\right|}\ .$ (8)
This approach drastically reduces the overall computational cost by increasing the accuracy of the long-range contributions, allowing the selection of smaller short-range regions whose interactions are substantially more expensive to compute. In particular, with a proper selection of the short-range cutoff, which dictates how many MM atoms are treated in short- and long-range domains, and the order for the multipole expansion, it is possible to reach the same accuracy as if all the electrostatic interactions were treated exactly, i.e. according to Eq. III.1, yet, at a fraction of the cost, as shown in Fig. 4.
A detailed theoretical treatment of the electrostatic embedding QM/MM scheme implemented using MiMiC can be found in Ref. 34. It includes explicit expressions for the polarizing potential acting on the QM subsystem due to MM atoms, i.e., the functional derivative of the QM/MM interaction energy with respect to the electron density, and for the forces acting on QM and MM atoms, i.e., the derivatives of the QM/MM interaction energy with respect to the atomic positions.
@@
Electrostatic embedding QM/MM is implemented using MiMiC by interfacing two external programs, a QM and an MM client, to an MD driver. The general simulation workflow is illustrated in Fig. 5. After the initialization, consisting of data collection and allocation, the simulation enters the main MD loop. At each MD step, MiMiC distributes atomic coordinates among the clients, which then concurrently compute energy and forces that are collected by MiMiC and passed to the MD driver to update velocities and subsequently calculate a new set of coordinates.
@@
The QM client solves the QM equations in the presence of the external MM potential computed by MiMiC, sends the converged electron density to MiMiC, and then proceeds to compute the QM forces while MiMiC computes the corresponding electrostatic QM/MM contributions. Finally, MiMiC collects energies and forces from all client programs and forwards them to the MD driver, which then updates velocities and positions starting a new MD step.
@@ III.2 Polarizable Embedding QM/MM MD
In electrostatic embedding, the QM system is polarized by the electric field generated by the fixed point charges of the MM subsystem. As such, this approach does not account for the explicit polarization of the MM subsystem, which is important in many cases and even crucial in some, e.g., in processes involving absorption or emission of photons and electron or proton transferBondanza et al. (2020). Polarizable embedding QM/MM schemes address this issue by using an MM force field that includes a polarizable component. For instance, in the frequently used induced point dipole model, the additional dynamic degrees of freedom are modeled using electric polarizabilities that describe an electronic charge distribution distorted by an electric field. Here, we briefly summarize the theoretical basis for polarizable embedding QM/MM MD using the polarizable AMOEBAPonder et al. (2010) force field to model the MM subsystem (QM/AMOEBA). For further details, we refer to Refs. 47; 48.
The electrostatic interactions in the AMOEBA force field are modeled using permanent multipoles, up to and including quadrupoles, and induced dipoles, from which electrostatic and polarization QM/MM interaction energies arise. The former is computed similarly as in electrostatic embedding QM/MM using modified interaction tensors (Eq. 5). The polarization energy is expressed in terms of induced dipoles and electric fields, so it is necessary to first determine the induced dipoles through coupled polarization equations. A characteristic aspect of AMOEBA is its non-variational nature, which stems from the fact that the electric field from the permanent multipoles used to determine the induced dipoles differs from the one used to calculate the polarization energy. The former is referred to as the direct field and the latter as the polarization field. The difference appears due to the use of different scaling factors and exclusion rules for each of these fields. Consequently, it is not possible to derive the polarization equations by simply requiring the stationarity of the energy functional with respect to the polarizable degrees of freedom. To avoid the complexity of non-variational optimization, a variational formulation has been adopted, resulting in two sets of polarization equations and, therefore, two sets of induced dipolesLoco et al. (2017); Nottoli and Lipparini (2020).
@@
Enabling fully polarizable QM/MM MD simulations using MiMiC requires modifications of the algorithm used to solve the QM equations. Achieving full self-consistency entails iteratively adjusting the induced dipoles in the MM subsystem and the electron density in the QM subsystem until both are converged with respect to the field generated by the other subsystem. This iterative process is characteristic of most polarizable embedding QM/MM models. The general workflow of a polarizable embedding QM/MM MD simulation using MiMiC is quite similar to the one for electrostatic embedding QM/MM shown in Fig. 5. It requires extending the Solve QM equations step with the iterative cycle illustrated in Fig. 6.
@@ III.3 Multiple Time Step QM/MM MD
MTS algorithms reduce the cost of MD simulations by separating forces into fast and slow components and integrating the latter less frequently, i.e., with a larger time step. Assuming a time-scale separation between these components, it is possible to split the classical Liouville operator of an $N$ particle system as
@@
Figure 7: Schematic workflow for MTS QM/MM MD using MiMiC.
MTS-generated trajectories are, by construction, at the high level of theory with the outer time step $\Delta t$, regardless of the low-level method used. The choice of the latter only determines the magnitude of the correction and, hence, the oscillations of the slow force component over time. These oscillations dictate the maximal ratio between the inner and outer steps that can be applied to generate stable dynamics.
@@
Performing a full MTS MD step requires executing $n$ inner steps at a lower level of theory (as shown in Eq. 12) followed by a single high-level outer step. Fig. 7 depicts a schematic workflow of an MTS QM/MM MD algorithm using MiMiC. At every inner step, MiMiC collects the low-level forces from the QM and MM clients and computes the low-level QM/MM force contributions (Eq. 14a), which together form the fast forces (Eq. 13a). If the current step is not a multiple of $n$, velocities and coordinates are updated at the low level with the time step $\Delta t/n$; else, at every $n$-th step, the outer loop is also undertaken. In the outer loop, MiMiC additionally collects the high-level QM forces, computes the high-level QM/MM force contributions (Eq. 14b), and then proceeds to calculate the slow forces (Eq. 13b). These are used by the MD driver to update the velocities, thus completing a full MD cycle with the time step $\Delta t$ at the high level of theory. Note that with MiMiC, it is straightforward to use two different QM client programs for the low- and high-level forces.
IV Features
The first development version of the MiMiC framework was released in 2022 on the occasion of the CECAM Flagship School Multiscale Molecular Dynamics with MiMiCmim (2022) held at the CECAM Headquarters at EPFL in Lausanne, Switzerland. It implements the DFT-based electrostatic embedding QM/MM MD method illustrated in Sec. III.1, using GROMACS Abraham et al. (2015); Páll et al. (2015) as the MM client and CPMD cpm (2023) as the QM client and MD driver. Furthermore, it also supports MTS QM/MM MD (Sec. III.3) through CPMD, and, through the PLUMED library, QM/MM-based enhanced sampling simulations and free energy calculationsBonomi et al. (2009); Tribello et al. (2014); Colón-Ramos et al. (2019).
@@ IV Features
Our initial choice for an MM client was GROMACSPáll et al. (2015), which is a free and open-source software suite for classical MD, distributed under LGPL version 2.1, whose developers aimed to provide the highest possible performance and efficiency on any hardwareAbraham et al. (2015). It implements a native heterogeneous parallelization setup using both CPUs and GPUs with the possibility to offload all force components to GPUs in single-precision calculations. The latest versionsPáll et al. (2020) also introduced new direct GPU–GPU communication and extended GPU integration, enabling excellent performance across multiple GPUs and efficient multi-node parallelization. Moreover, GROMACS supports a large variety of popular biomolecular force fields, including AMBERPonder and Case (2003), CHARMMBrooks et al. (2009), GROMOSScott et al. (1999), and OPLSRobertson, Tirado-Rives, and Jorgensen (2015), and provides a wide array of well-documented pre- and post-processing tools. New users can also benefit from an extensive range of tutorials, documentation, and an active user community.
Figure 8: Parallel scalability of a MiMiC-based QM/MM implementation using GROMACS as the MM client and CPMD as the QM client and MD driver. The system comprised 130 828 atoms, of which 142 were treated at the QM level using the hybrid B3LYP functional. The speed-up is calculated based on the CPU time required for one MD step, normalized to a reference run using seven nodes. Data from Raghavan et al.Raghavan et al. (2023b) who ran the simulations on the CPU partition of the JUWELS cluster at the Jülich Supercomputing Centre.
@@
V Upcoming Features and Client Programs
@@
These efforts include the implementation of dynamically generated restrained electrostatic potential (D-RESP) chargesLaio, VandeVondele, and Rothlisberger (2002b), extending the original approach to also fit higher-order atom-centered multipoles. Such an extension will also make it possible to implement different coupling schemes for electrostatic interactions, for example, by introducing a three-level approachLaio et al. (2004), which would significantly increase electrostatic embedding efficiency, particularly for elongated QM subsystems. In addition, the implementation of D-RESP charges will also serve as an on-the-fly analysis tool for tracking ongoing changes in the electronic structure and will also enable charge-driven enhanced samplingSulpizi, Rothlisberger, and Laio (2005). Another important addition, building on D-RESP, will be the development of QM/MM force-matching schemes, such as those in Refs. 91 and 92, in which the total forces on the QM atoms, as well as the QM electrostatic potential and field, are extracted during a QM/MM MD simulation, for a set of sampled configurations. These quantities are then used to parameterize standard or polarizable biophysical force fields for the QM subsystem. Such QM-matched force fields can be employed to achieve more extensive sampling in a fully classical MD simulation of the whole system.
@@
A different extension of capabilities in MiMiC will be based on enhancing the MTS workflow illustrated in Sec. III.3, which unlocks a plethora of possible combinations of different high- and low-level methods, including wave function, DFT, semiempirical, force field, and ML modelsMouvet et al. (2022). A future release of MiMiC will also include an adaptive ML-accelerated MTS QM/MM scheme in which the ML model is retrained on the fly whenever the system accesses new regions of configurational space (e.g., during a chemical reaction) where the original ML model has insufficient predictive power. In addition, this adaptive scheme allows for a dynamic training set build-up during the run, thus circumventing the need for extensive and computationally demanding a priori generation of high-level reference data.
@@
Regarding the development of new interfaces to external client programs, we currently focus on achieving higher performance and higher accuracy for both QM and MM subsystems. Thus, we strive to enable large-scale QM/MM MD simulations on current and future (pre-)exascale supercomputers by coupling to highly scalable GPU-enabled QM client programs, namely, CP2K Kühne et al. (2020), DFT-FE Motamarri et al. (2020); Das et al. (2022), and Quantum ESPRESSO Giannozzi et al. (2009); Carnimeo et al. (2023), as well as making available more accurate wave-function methods, as implemented in CP2K and CFOUR Matthews et al. (2020); Stanton et al. . As a further step towards higher accuracy, we are implementing a fully polarizable QM/MM method (see Sec. III.2). The pilot implementation will use CPMD as the MD driver and QM client and the high-performance Tinker-HP Lagardère et al. (2018); Adjoua et al. (2021) package as the MM client. However, the implementation in MiMiC will be general, and support for other clients, such as OpenMM Eastman et al. (2012, 2017, 2023), will follow soon after.
@@
To achieve these goals, work on interfaces with several client programs is currently underway. In the remainder of this section, we review their distinctive features, which make them ideal candidates for integration into the framework. As new programs will be coupled to MiMiC, the pre-processing toolkit MiMiCPy will also be augmented accordingly.
@@
CP2KKühne et al. (2020) is a free and open-source software package, distributed under the GNU General Public License (GPL) version 2.0 or later, for simulations of solid-state and (bio)molecular systems, giving access to a broad range of electronic structure methods. It focuses on ensuring excellent performance, thanks to algorithms developed for modern HPC systems, also including GPU support. In particular, the electronic structure module QuickstepKrack and Parrinello (2004); VandeVondele et al. (2005); Kühne et al. (2020) provides an efficient infrastructure of integral routines and optimization algorithms for a wide spectrum of methods, including orbital-free DFT, KS-DFT, and MP2. Quickstep is based on mixed Gaussian and PW (GPW) basis sets and their augmented extension (GAPW), where auxiliary PW basis sets are used within a GTO scheme. Recently, CP2K’s linear scaling DFT implementation was used for calculations on the SARS-CoV-2 spike proteins with almost 83 million atoms, for which it scaled up to 4 400 NVIDIA A100 GPUs with a parallel efficiency of over 80%Schade et al. (2023).
@@
DFT-FE Motamarri et al. (2020); Das et al. (2022) is a recently developed free and open-source code, distributed under LGPL version 2.1 or later, that implements real-space KS-DFT based on a spatially adaptive and systematically convergent higher-order spectral finite-element (FE) basis. It accommodates both norm-conserving pseudopotentials and all-electron calculations, with ongoing implementation efforts to support ultrasoft projector-augmented-wave pseudopotentials. DFT-FE enables fast and accurate large-scale DFT calculations on generic materials systems involving many tens of thousands of electrons on both many-core CPU and hybrid CPU-GPU architectures. The good parallel scalability and efficient use of GPUs have enabled systematically convergent DFT calculations with low wall times, e.g., an ab initio MD step of systems containing 10 000-20 000 electrons can be completed in wall times of <1 minute Das et al. (2022). In addition to the support of NVIDIA GPUs through CUDA, DFT-FE’s GPU porting layer was recently extended to support AMD GPUs through HIP, with its exascale capability recently demonstrated through scaling up to 8 000 GPU nodes (32 000 AMD MI250X GPUs) of the Oak Ridge Leadership Computing Facility Frontier exascale supercomputer Das et al. (2023).
@@
Quantum ESPRESSOGiannozzi et al. (2020) is an integrated and free suite of programs, distributed under GPL version 2.0 or later, designed for first-principles electronic structure calculations and materials modeling at the nanoscale. In recent years, it has been refactored and reorganized in multiple layers of codeCarnimeo et al. (2023), consisting of loosely coupled modules and libraries, which are tailored to be independently extensible and maintainable. Recent versions consist of a core computational layer for DFT simulations based on PWs and pseudopotentials, containing a quantum-engine package to solve self-consistent KS equations for periodic systems (PWscf). To achieve high performance on different architectures, Quantum ESPRESSO features a layered, fine-tuneable parallelization scheme based on MPI and OpenMP, as well as GPU support. Porting to GPU has been achieved through CUDA Fortran and, to an increasing degree, directive-based programming models (OpenACC and OpenMP) for offloading to heterogeneous platforms, making it as architecture-agnostic as possible. The optimal combination of all available parallelism levels currently allows the Quantum ESPRESSO suite to achieve a substantial increase in performance and efficiencyCarnimeo et al. (2023).